According to the Census Bureau, in October 2016, the average house price in the United States was $27,458. 5 years earlier, the average price was $22,508. What was the annual increase in the price of the average house sold?

Multiple Choice

  • 3.65%

  • 4.87%

  • -3.90%

  • 4.46%

  • 4.06%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

27,458=22508*(1+r/100)^5

(27,458/22508)^(1/5)=(1+r/100)

(1+r/100)=1.0406

r=1.0406-1

=4.06%(Approx).